OpenStudy (anonymous):

What is the slope of the line defined by the equation 8x + 2y = 5?

OpenStudy (cwrw238):

rearrange to the for y = mx + c where m = slope 2y = -8x + 5 divide by 2: y = -4x + 2.5 - now you have the slope

OpenStudy (cwrw238):

compare these 2 equations and you'll see what the slope is y = m x + c y = -4 x + 2.5

OpenStudy (anonymous):

How do I do that?

OpenStudy (cwrw238):

m = slope what in the second equation coincides with m?

OpenStudy (cwrw238):

whats directly under the m in second equation? - what precedes the x?